Circular No. 8815

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2002 GZ_{31}

K. S. Noll, Space Telescope Science Institute (STScI); W. M. Grundy, Lowell Observatory; S. D. Kern, STScI; H. F. Levison, Southwest Research Institute; and D. C. Stephens, Brigham Young University, report the detection of a binary companion to the transneptunian object 2002 GZ_{31} (cf. MPECs 2002-K12, 2003-M16; MPO 66276). The observations were made during 2006 Apr. 23.352–23.377 UT with the High Resolution Camera of the Advanced Camera for Surveys on the Hubble Space Telescope, using the clear filters with one 300-s exposure at each of four dithered positions on the detector. The two components were separated by an angular distance of 0′′.070 ± 0′′.009 and differ in brightness by 1.0 magnitude. The fainter component lies at a position angle of 352^{o} ± 10^{o} from the primary. The projected separation of the objects in the sky plane is 2060 ± 270 km.

COMETS C/2006 X3–X10 (SOHO) Additional comets have been found on SOHO website images (cf. IAUC

8813) — all Kreutz sungrazers except for C/2006 X10 (Meyer group). K. Battams notes that C/2006 X3, C/2006 X5, C/2006 X6, C/2006 X8, and C/2006 X9 were all extremely faint (mag ∼ 8.5). C/2006 X3, which was observed during a solar-particle storm that hit the SOHO spacecraft, appeared diffuse. C/2006 X4 was stellar in appearance and reached mag ∼ 6 in C3 images, while C2 images showed a hint of a very faint tail. C/2006 X5 was very diffuse and elongated. C/2006 X6 was tiny and slightly diffuse. C/2006 X7 was stellar in appearance in C3 images, of mag ∼ 6.5, while C2 images showed it as slightly diffuse with a hint of a faint tail. C/2006 X8 was stellar in appearance. C/2006 X9, which was also found by R. Kracht, was a very diffuse blob. C/2006 X10 was stellar in appearance, reaching mag ∼ 7.5.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 X3 Dec. 9.567 17^{h}06.^{m}8 −24^{o}43′ C2  BZ 2007-C40
C/2006 X4  10.513 17 08.9 −26 51 C3/2 TS 2007-C40
C/2006 X5  10.550 17 11.4 −24 49 C2  MM 2007-C40
C/2006 X6  11.675 17 16.4 −24 53 C2  XG 2007-C40
C/2006 X7  11.988 17 17.0 −26 10 C3/2 HS 2007-C41
C/2006 X8  12.567 17 20.9 −24 55 C2  BZ 2007-C41
C/2006 X9  15.204 17 34.6 −25 01 C2  TH 2007-C41
C/2006 X10  15.479 17 30.3 −22 17 C2  RK 2007-C41
2007 February 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green